Usse the variable expression 3x - 1
Evaluate the expression when x= -3, -2, -1, 0, 1, 2, and 3

Answer by checkley77(12569)

3x-1
3*-3-1=-9-1=-10 answer.
3*-2-1=-6-1=-7 answer.
3*-1-1=-3-1=-4 answer.
3*0-1=-1 answer.
3*1-1=3-1=2 answer.
3*2-1=6-1=5 answer.
3*3-1=9-1=8 answer.